A circuit of 500 V for single-phase asynchronous motor has synchronous impedance of 3.2 Ω. The armature resistance is 0.3 Ω. Determine the induced emf if the machine develops 31.3 kW and the mechanical losses equals 4 kW


1
Expert's answer
2021-08-30T01:54:15-0400

"E= \sqrt{(V \cos \phi+IR)^2+(V \sin \phi+IX)^2}\\
E= \sqrt{(\frac{500}{\sqrt3}+IR)^2+(IX)^2}\\"

Input power

"4 kW +31.3 kW\\
35.3kW"

"35.3kW= \sqrt3 VI \cos \phi\\
I= \frac{35.3*10^3}{\sqrt3 *500}*40.76A\\
\implies E_{phase}= 327.94 V"